Sp2 in the first graph. 13 and 1

  

Sp2 in the first graph.

   Sp2 in the first graph. 13 and 1

13 and 14

   Sp2 in the first graph. 13 and 1


   Sp2 in the first graph. 13 and 1
Answer #1


    average acceleration between 4-8 s is area of square (4*4+24)= 40 /2 = 20 m/s2

13. initial velocity u = 14 m/s, acceleration a = 2.5 m/s2, t= 3 s
   a. velocity at the end of the 3 s is using equations of motion


   v = u +at = 14+2.5*3 = 21.5 m/s

   b. distance covered during the time

            s = ut + 1/2 at^2
            = 14*3+1/2*2.5*9
       = 53.25 m
14.
   initial velocity u = 2 m/s, acceleration a = 2.2 m/s2, t= 3 s
   a. velocity at the end of the 3 s is using equations of motion


   v = u +at = 2+2.2*3 = 8.6 m/s

   b. distance covered by the runner during 3 s is


   s = ut + 1/2 at^2
            = 2*3+1/2*1.1*9
       = 10.95 m